Three charges are fixed in the x?y plane as follows: 1.5 nC at the origin (0, 0); 2.4 nC at (0.75m , 0); –1.9 nC nC at (0, 1.50 m ).

Part A

Find the force acting on the charge at the origin.

Express your answers using two significant figures. Enter the x and y components of the force separated by a comma.PHYS203 - SPRING I 2016 - RIZVANOV Signed in as joshua emswiler Help Close Homework1a Problem 15.46 Resources 《 previous | 4 of 6 | next Problem 15.46 Part A Three charges are fixed in the -y plane as follows: 1.5 nC at the origin (0, 0);2.4 nC at (0.75 m,0),-1.9 nC nC at (0, 1.50 m). Find the force acting on the charge at the origin Express your answers using two significant figures. Enter the x and y components of the force separated by a comma. F. F Submit My Answers Give Up Incorrect, Try Again; 4 attempts remaining Provide Feedback Continue
Answer #1

F = k q1 q2 / r2

F(x) = -1.5*2.4*9*10-9*10-18 / 0.752 = -5.76*10-8

F(y) = 1.5*1.9*9*10-9*10-18 / 1.52 = 1.14*10-8